WHT diary

This is an informal log of noteworthy events in the life of the WHT. Sources of more detailed information include:

1987
----
Oct    Primary aluminised?


1988
----
Nov    Primary aluminised, secondary aluminised (twice)


1989 (observing with TAURUS, FOS, GHRIL, FLEX, QUBES)
----
May 10  One-month standdown to commission Cass A&G box.
Jul 10  6-week standdown to commission ISIS.
Sep  4  1-month standdown for dome.
Nov  7  Primary aluminised.
Dec 14  ISIS polarisation commissioning begins.


1990
----
Jan  3  MES (Manchester Echelle) run.
Feb  7  CFX (control-file observations) commissioning - never adopted.
Mar  5  Secondary standdown.
Apr 16  Autofib (1?) on telescope.
Jun  4  1-month Cass rotator standdown.
Oct  1  1-week dome-shutters standdown.
Nov 28  Primary aluminised


1991
----
Jun 24  UES commissioning begins.
Jul 19  First FAST run (Genzel).
Nov 18  Primary aluminised.


1992
----
Mar 16  LDSS commissioning begins.
Jun 12  First Leicester-photometer run.
Oct  6  Primary aluminised.
Nov  2  2-week tube-straightening standdown.


1993
----
Jan     Martin Fisher measured differential flexure between top and bottom
        ends -0.56 +- 0.24 mm, confirming improvement due to starightening
        and small enough to allow PF commissioning to go ahead.
Jan     Ladder installed to allow safer access to cube from UES platform.
Jan     Top end of telescope being cleaned.
Jan     Guide-star server installed?  Reported in Gemini 3/93.
Feb     Cable wrap amplifier burnt out.
Feb     One of dome motor brakes burnt out, dome filled with fumes.
Feb     Rudi Le Poole carrying out thermal audit.
Mar     8 large instrument-storage cupboards installed in gantry.
Mar     New 6100-A dichroic and longer BFOLD for ISIS installed by SPW.
Mar     Duckboards installed outside WHT to leave ice-free path.
Mar 19  Ethernet shorted in cable wrap, now using spare.
Mar 29  Standdown until Apr 1 to replace cable wrap.
        Inspection of cable wrap revealed many twisted and torn cables.
        Cables and links replaced.  
        WYFFOS table and supports installed during standdown.
Mar 30  INT mirror and 2 IAC mirrors aluminised in WHT tank.
Apr     Old optics lab emptied of equipment in preparation for conversion to
        instrument-test area.  Aluminising area also emptied of instruments
        and cupboards.
Apr     Tea room refurbished (double in size, new units)
Apr     Daylight IR observing of SN1993J by team from IAC (starting half
        an hour before sunset).
Apr     Extensive observations of SN1993J with UES.
May     Welder visited from Tenerife to mend cracked welds in WHT dome rail.
Jun  1  Standdown until Jun 3 to inspect new cable wrap and replace cables.
Jun 12  Hitchhiker team observing until Jun 26 (last run this year).
Jun 27  Open day.
Jun 30  Leap second inserted.
Jul 18  Open day.
Aug     Microthermal sensor array installed by Brian Hucklesby.
Aug     Imaging polarimetry commissioned.
Aug  4  NOT mirror aluminised in WHT tank.
Aug 15  Open day.
Aug 29  Open day.
Sep  5  Open day.
Sep 12  Open day, last of year.  4130 visitors on the 6 open days.
Sep 15  Crane tests.
Oct 12  Dave Gellatly installed PF electronics crates and fibre trunking.
Oct     Large numbers of tourists leaking onto observatory site because
        Roque road closed.
Oct 21  Primary aluminised - Aluminising standdown until Oct 24.  
Nov 22  PF commissioning until Dec 5.
Dec     Martin Fisher found that movement of images at GHRIL was due to
        Nasmyth mirror seating badly against a bracket.  Realigned.
Dec 24  Run with ARNICA IR instrument until Jan 2.


1994
----
Jan     Worst storm for 5 years left domes (and everything else on mountain)
        thickly sheathed in ice.  Some minor damage from falling ice.
Jan 20  PF commissioning until Jan 31, nearly wiped out by bad weather.
Feb     Ray D showed video taken by Martin Fisher from prime focus, showing
        hot air billowing across altitude bearings on either side.
Feb     Control-room partitioning installed.
Mar     Schedule in middle of month changed around to accommodate 
        Vilchez (he had to observe in Hawaii).  Net effect is that
        2 dark service nights become slightly brighter.
Mar     ISIS pointing 1.5", UES 3"
Mar  1  Angela Baldwin (great-great-granddaughter of William Herschel) and
        husband visited WHT, and presented us with portraits of William and
        John Herschel.
Mar  3  Manual logging of observations stopped; automatic logging
        begun (LOGFORMAT format 12).
Mar 12  Dave King installed new dichroics 4500, 5400, 5700, 6100, 7500
        on 5 mounts, each with 1 dichroic, 1 mirror.  Old dichroics retired.
Mar 15  New door being installed between control room and corridor.
Apr     Prime-focus commissioning completed.
May  2  GHRIL autoguiding commissioned.
Jun  1  WHT active at 3 focal stations in one night (ISIS, UES, GHRIL).
Jun 14  Crosswires removed from tube because diffraction spikes interfering
        with coronograph observations of disks round stars.
Apr 15  MPE's 3D IR camera on until Apr 21
Apr 22  PF commissioning until May 31 Jul 18  Moreno observing with Coronograph
Sep 19  UES A&G upgraded, bigger fibres, better throughput
Oct 11  AUTOFIB2 commissioning until Oct 20
Nov 14  3-night standdown for aluminising, replacement of data optical fibres
        by 62.5-micron industry-standard fibres, move of electronics racks
        to back of control room
Nov 15  Shack-Hartmann at several elevations

1995
----
Jan 25  GHRIL IR-derotator chassis aligned by Sue Worswick.  Derotator will
        be aligned with chassis in Cambridge.
Mar 10  New control desk installed by John Mills, Clive Amos et al.
Mar 30  Ken Pounds (PPARC chief exec) visits ING
Jun     ISIS optics serviced (SPW/DLK)
Jul 10  David Walker's UCL team testing ISAAC adjustable collimator on
        telescope
Oct  3  Observed all night with mirror support off (by accident!)
Oct 4-7 CO2-snow cleaning tests (reflectivity increased 80% to 85%) and
        aluminising of primary.
Oct 8   Shack-Hartmann test data rubbish because wrong spacer in
Nov 20  Secondary mirror aluminised during prime-focus run.
Nov     LORAL CCD commissioning cancelled; 6 half-nights released for
        S/D work.
Nov 24  Secondary transducer no.1 dead (dead at least since Aug 14).
        Fixed next day.
Dec 23  Cass TF term set to -10

1996
----
Jan 10  New (Nov 95) Cass pointing model entered by JMB.  Rms now 1.1"
        during CALIBRATE
Jan 11  Primary, secondary and Nas flat CO2-snow-cleaned by Mariet.  Nas flat
        showed some improvement
Jan 16  Nasmyth flat cleaned with PL25 spray-on, peel-off, noticeable
        improvement (less grainy appearance)
Feb 20  Nobel-prize winning medics visit WHT, but only have 5 mins, because
        they have to rush for a flight to Lanzarote
Mar 18  Shack-Hartmann tests at aux port (CRB/SPW) show sig more spherical
        aberration than autumn 1994, probably consistent with numbers from
        aux-port and TV out-of-focus imaging
Mar 18  Newly-coated IR derotator being installed at GHRIL (SPW/MBROX)
Mar 24  Nasmyth flat removed
Mar 25  Nasmyth flat aluminised (first time for years), reinstalled
        Mar 27 after slight delay due to failure of glue on one pad
Apr  3  5000x8000 NAOJ mosaic camera goes on PF for 2 weeks
Jun 27  Nasmyth mirror cleaned with PL25
Jun 29  Royal opening of Galileo; visitors see Galileo mirror in
        WHT aluminising area
Jul 30  Secondary (primary according to MBROX) snow-cleaned.
Sep 30  Primary (only) aluminised, secondary looking good.
        No reflectivity measurement (new monitor acquired at time of 
        aluminising).


1997
----
May  5  Nas flat PL25-cleaned
Jun     Nas flat misaligned??
Jul  6  Nas flat PL25-cleaned (refl. .85 before, .86 after) 
Oct 15  TCS commissioning cancelled (software group not ready), 
        reverted to S/D + INTEGRAL
Oct 29-30  Site-properties conference in Los Cancajos
Oct 30  Prins Willem (NL) visits at night, opens Dutch open-tower solar 
        telescope Oct 31 and visits WHT dome during day
Nov  9  Major CCD electronics / DICI problem, both arms of ISIS unusable,
        5.5 hours lost (largest downtime for years)
Nov 10  Aluminising primary mirror.


1998
----
Jan  8  CIRSI run.
Jan 10  Two observers (WHT and INT) taken ill with flu?, whisked off site
Mar     UES slit jaws not going to reported width.  Traced in early Jun
        to dekker scraping against jaws.
May  22 Final run of Japanese mosaic CCD at WHT PF.
Jun  2  4 of 6 Nasmyth-flat pads being reglued off telescope (other 2 
        were reglued last (?) year).  2 of the 4 came loose ver easily,
        so flat may not have been positioning correctly.
Jun  3  Nasmyth flat back in, awaiting alignment
Jun  9 - 11  JOSE run, incl one night data in good seeing with simultaneous
        DIMM observations.  They found the IR derotator throughput to be
        poor (dirty surfaces).
Jun 15  Secondary baffle off (for K skybr measurements)
Jun     Az oscillations, +- 0.5 arcsec, about 2 Hz, after WHIRCAM put on
        at Cass?
Jun 23  Rotator collimation correction being incorrectly calculated by
        new TCS - claim for paybacks possible
Jun 27  Long-slit UES derotator not installed for Vors run - claim for
        payback possible.
Sep  7  Aluminising primary mirror.
Sep 10  Shack-Hartmann tests at Cass, FWHM < 0.2 arcsec
Sep 11 - Oct 1  
        CALIBRATEs ~ 5 arcsec.  Time service discovered to be 1 sec out.
Oct 12  Cass TT moved 5 deg out of synch with cable wrap, trashed a
        linear potentiometer.  No damage in wrap.
Oct 27  Final run of WHIRCAM as common-user instrument.


1999
----
Jan 17? Opening WHT after storm, buckets of water poured out of A&G box.
Jan 17  2.5 hours lost to lightning damage (of CAMAC), and 2 hours more
        Jan 23 (failure of UES rotator encoder).
Jan 20  Telescope driven accidentally into a dewar, damage to one clamp
        on a cryostat.
Feb  1  SAURON commissioning begins.
Feb  2  SCAM commissioning begins.
Mar  3  Final run of TAURUS as common-user instrument.
Apr~10  A&G mirrors all PL25 cleaned (BaRT), need alum after the flood
May  6  1 kg pulley-weight fell onto bare mirror from PF.
        Some permanent scratches.  1 night lost.
        CALIBRATE coefficients same on nights before and after, so
        mirror not moved.
May 13  PF observers report elliptical images (ell ~ 0.05 - 0.1),
        in elevation direction.  
May 26  ELECTRA commissioning, hampered by having to use WHIRCAM as
        detector.
Jun 21  ELECTRA commissioning.
Jul  1  CIRSI run.
Aug 25  Power cut early morning for half an hour, backup batteries
        failed, all day needed to recover systems.
Sep 21  TEIFU commissioning.
Oct     PF observer working at dec = -4, images marred by dozens of tracks
        of geostationary satellites (this turns out to be the dec at which
        they lie when viewed from the latitude of La Palma).
Oct  8  SAURON run.
Nov  1  TRIFFID run.
Dec  8  SCAM commissioning.
Dec 31  WHT scheduled closed (first night ever).


2000
----
Jan  1  No time lost to Y2K bugs.
Feb  7  Thomas Augusteijn deputy WHT telescope manager from today.
Feb 10-16  A&G and ISIS off.  INGRID folding flat installed in A&G.
        ISIS red-arm focus problems investigated.  ISIS slit cleaned.
Mar  2  Final LDSS run as common-user instrument.
Mar 16  INGRID commissioning.
Mar 27  SAURON run.
Apr 24  200th anniversary William Herschel's discovery of IR radiation
Apr 25  SCAM commissioning delayed by delay in delivery of liquid He.
Jun  1  Prince Felipe of Spain visits, takes an image of M51, files
        no defect reports.
Jun  2  Seeing 0.8 arcsec at all 3 telescopes at start of night,
        1.0 arcsec at all 3 telescopes at midnight.  First recorded
        occasion when ~ no dome seeing in any of the 3 telescopes?
Jun 13  May 1999 gouges in primary mirror ground out to relieve stress.
        Two patches of surface ~ 1 cm in diameter have been lost.
        Mirror trial washed with water (5% improvement in reflectivity)
        and then with water/detergent (1.5% further improvement) before
        aluminising.
Aug 14  NAOMI commissioning hampered by telescope oscillations,
        now almost eliminated by fine-tuning telescope-control feedback
        loop.
Sep  1  SAURON run.
Sep  9  NAOMI commissioning.
Sep 18  Stannddown night for installation of He pipes in Cass cable wrap.
Sep 29  SCAM run (first half night lost after failure of motor in cooler).
Dec 20  First NAOMI science run wiped out by bad weather.

winter 2000/01 - best Nov/Dec/Jan/Feb weather for at least 10 years


2001
----
Jan     ISIS data-acquisition switched to ultradas.
Mar  6  Commissioning INGRID with new aspheric.
Mar 14  SAURON run.
May 29-31, Jun 4-15 
        NAOMI commissioning + 2 nights service observing
Jun 26  Bent gripper found in AF2, replaced.
Sep 23  Severe coma introduced after focus drive sent to limit
        (after replacement of drive belt, which came off), 
        causing a counterweight to crush a cable connector and
        tilt the secondary mirror 1 mm off its kinematic mount.
Oct 17  Bottom shutter repaired.
Oct 17  ING board visit WHT.


2002
----
Jan 26  Start of commissioning run for Rayleigh laser guide star.
Jan 30  FOS finally removed from ISIS (improves balancing).
?       UES removed from Nasmyth platform
Dec     GRACE dedicated AO enclosure installed on old UES paltform

2003
----
Mar     NAOMI moved into GRACE
Oct 8   ING board visit WHT

2005
----
Sep     Bush fires
Sep 28  Service night with ISIS red and blue, aux, INGRID and OASIS - record?

2006
----
Jun 11  Spanish minister of science and educ (MEC) visits WHT, sees NAOMI
Aug     (or Sep) rain fell on WHT, date unknown, rainspots cover ~ 10% of
        surface
Oct  6  Half-metre area of primary experimentally washed (half with soap)



Chris Benn (crb@ing.iac.es)

Last updated 2006 Oct 4

Click here to return to WHT home page